OpenStudy (anonymous):

i know it's she sells 24 item.., ok??? (700 - 100)/25 = 24 how would u conclusion as a complete sentence

OpenStudy (mayankdevnani):

600/25=120/5=24

OpenStudy (mayankdevnani):

yaa you're right>.

OpenStudy (mayankdevnani):

@dietrich_harmon

OpenStudy (anonymous):

how would u conclusion as a complete sentence

OpenStudy (mayankdevnani):

but where is your sentence

OpenStudy (mayankdevnani):

or question

OpenStudy (mayankdevnani):

Sharon earns $25 per item she sells plus a base salary of $100 per week. Write and solve an inequality to find how many items she must sell to earn at least $700 per week. Provide your conclusion as a sentence.

OpenStudy (mayankdevnani):

is it your question

OpenStudy (mayankdevnani):

conclusion=so he sells 24 items per week

OpenStudy (mayankdevnani):

@dietrich_harmon

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

thank u sweetie